# Baltimore Atlas: FreqWeaver Adapter for Semi-supervised Ultra-high Spatial Resolution Land Cover Classification

### v1

@@ 4.4. Comparison With State-of-the-Art PEFT Methods
To evaluate the effectiveness of the proposed adapter, we conducted extensive comparative experiments against several state-of-the-art parameter-efficient fine-tuning (PEFT) methods, including the SAM Adapter(Wu et al., 2023b), LoRA(Hu and Shen, 2022), Conv-LoRA(Zhong et al., 2024), and the Earth Adapter(Hu et al., 2025). These approaches represent diverse strategies for adapting large-scale models like SAM2, with different focuses on domain specificity and parameter efficiency, as summarized in Table 1.
@@
Limitations of LoRA and Conv-LoRA. Our experimental results show that both LoRA and Conv-LoRA achieve competitive performance with remarkably low parameter overhead. LoRA, which leverages low-rank decomposition for efficient adaptation, is widely adopted for lightweight fine-tuning tasks. However, due to its linear structure, it exhibits limited capability in modeling the spatial complexity of high-resolution remote sensing scenes, which often involve multi-scale objects, intricate semantics, and fine-grained textures. To address this limitation, Conv-LoRA enhances the original LoRA design by incorporating convolutional layers to strengthen local feature extraction. This modification improves its ability to model small objects and boundary details, leading to noticeable performance gains. Nonetheless, despite these improvements, Conv-LoRA still lacks the representational capacity to fully exploit the potential of high-resolution remote sensing imagery, particularly in capturing complex spatial-frequency interactions, leaving room for further enhancement.
@@
Limitations of Earth Adapter and SAM Adapter. The Earth Adapter, although specifically designed for remote sensing tasks, attempts to decouple frequency components without incorporating targeted spatial modeling mechanisms. This results in relatively high parameter consumption and unsatisfactory performance. The SAM Adapter, originally developed for medical image segmentation, shows moderate adaptability when transferred to remote sensing domains. However, due to the increased complexity and resolution of remote sensing data, it struggles to effectively model fine-grained structures and semantic variations.
@@
Effectiveness of the Proposed FreqWeaver Adapter. In contrast, our proposed FreqWeaver Adapter introduces a frequency-aware design tailored for remote sensing applications. By explicitly decoupling low- and high-frequency components and applying targeted processing strategies to each, FreqWeaver effectively suppresses low-frequency artifacts while preserving high-frequency structural details. This leads to improved segmentation accuracy and better adaptation to the SAM2 backbone. Importantly, FreqWeaver achieves this superior performance with fewer parameters than both the Earth Adapter and the SAM Adapter, demonstrating its strong domain-specific effectiveness and efficiency for high-resolution remote sensing tasks. We also present visualizations to validate the effectiveness of our method, as shown in 3.
@@
Table 2. Comparison with state-of-the-art Land Use and Land Cover (LULC) methods.
@@
4.5. Comparison with State-of-the-Art Land Cover Classification Methods
@@
To evaluate the effectiveness of our proposed approach, we conducted extensive comparative experiments against several representative state-of-the-art Land Cover and Land Use Classification (LULC) methods. Specifically, we selected five recent and competitive baselines from the literature (2021–2024)—ABCNet (Li et al., 2021), UNetFormer (Wang et al., 2022), CMTFNet (Wu et al., 2023a), RS3Mamba (Ma et al., 2024b), and SAM-RS (Ma et al., 2024a) chosen for their strong benchmark performance and representation of diverse architectural paradigms, including CNN-based, Transformer-based, and Mamba-based models. The quantitative comparison results are reported in Table 2, while qualitative visualization examples are provided in Fig. 5 to further illustrate the advantages of our method.
@@
Limitations of Existing Methods. Quantitative results (Table 2) demonstrate that the proposed method consistently outperforms all selected baseline methods. In comparison to the strongest baseline, SAM-RS, our model achieves an improvement of 3.44% in IoU and 2.03% in Dice score, underscoring its superior performance in LULC tasks. Despite the reasonable segmentation accuracy exhibited by prior methods, visual analysis reveals several limitations: (1) Misclassification remains prevalent in complex scenes, particularly those with high intra-class variance or subtle inter-class distinctions (e.g., differentiating various types of vegetation or urban structures); (2) Object boundaries are often rendered blurry and imprecise, a critical deficiency for ultra-high-resolution segmentation tasks that demand sharp delineations; (3) Models generally perform well on spectrally distinct classes like grass and trees, but falter with categories exhibiting complex geometries or semantic overlaps, such as distinguishing between parking lots and buildings, or between roads and barren land. Moreover, small object detection, such as identifying pools in residential areas, is often overlooked by existing approaches.
@@
Effectiveness of Our Method. To the best of our knowledge, our model is the first to adapt the Segment Anything v2 (SAM2) architecture for the task of LULC. By leveraging the rich prior knowledge encoded in this foundation model, we obtain robust category- and pixel-level representations. Building upon this foundation, the proposed FreqWeaver Adapter enhances the semantic granularity of feature modeling—by explicitly capturing and integrating multi-frequency information crucial for detail preservation—thereby enabling significantly improved delineation of object boundaries and detection of small structures. Notably, the enhanced feature representation from the FreqWeaver Adapter allows our model to effectively exploit contextual cues, such as the presence or absence of vehicles, to better distinguish between semantically similar regions like parking lots and roads. These capabilities demonstrate the strengths of our architecture in ultra-high-resolution LULC tasks. Additionally, the integration of a semi-supervised learning strategy further improves the robustness and generalization of our model under limited annotation scenarios.
@@
Figure 6. Visualization of Land Cover Classification.
4.6. Exploration of Downstream Applications in Map Analysis
To evaluate the practicality of our approach in real-world scenarios, we applied the proposed method to downstream ultra-high-resolution map analysis tasks, as summarized in Table 3. Despite being trained on a relatively small dataset, our method exhibits strong generalization capabilities, achieving an Overall Accuracy (OA) of 84.55% and a Kappa coefficient of 81.38%. These metrics reflect not only the correctness of predictions but also the statistical reliability of classification performance across diverse land cover categories.
A detailed breakdown of class-wise performance is provided in Table 3. Notably, high User’s Accuracy was achieved for categories such as Building (93.71%), Water (94.92%), and Grass/Shrub (91.80%), indicating precise identification of spatial regions with distinct spectral or structural characteristics. At the same time, balanced Producer’s Accuracy was maintained, particularly for Tree Canopy (98.98%) and Road (89.94%), demonstrating strong recall and comprehensive spatial coverage.
@@ 4.6. Exploration of Downstream Applications in Map Analysis
To further validate the effectiveness of our method, qualitative visualizations were performed on real-world map segments, as illustrated in Fig. 6. The results show that our approach successfully preserves object boundaries, captures fine-grained structures, and adapts effectively to spatial heterogeneity, confirming its suitability for downstream high-resolution map analysis tasks.
@@
4.7. Ablation Studies
To evaluate the contribution of each component within the proposed framework, we conducted extensive ablation studies, as shown in Table 4. In each experiment, specific modules—FreqWeaver (FW) and the Uncertainty-Aware Teacher Student Model (UATS)—were selectively enabled or disabled to analyze their individual impact on segmentation performance.
Effectiveness of FreqWeaver. We introduce FreqWeaver (FW) to decouple the frequency and temporal components within SAM2, enabling a parameter-efficient mechanism to capture complex patterns in remote sensing imagery. Experimental results show that, compared to full fine-tuning, our method achieves a 2.81% improvement in IoU and a 3.25% improvement in Dice score, while updating only 5.96% of the total parameters. These findings highlight the effectiveness of FreqWeaver in fine-tuning large vision models for remote sensing tasks and provide a scalable pathway for adapting foundation models under strict resource constraints.
@@ 4.8. Segment Anything 2 Adaptation
Segment Anything 2 is built upon the Hierarchical Detection Transformer (HieraDet), a multi-scale transformer architecture composed of four stages. In the original Segment Anything framework, adapters such as the SAM Adapter are inserted into every transformer block. However, directly applying this strategy to Segment Anything 2 often results in suboptimal fine-tuning performance due to its deeper and hierarchically structured design. To better accommodate these architectural differences, we conducted a series of empirical studies to investigate more effective adapter integration strategies. The results are presented in Table 5.
Specifically, we hypothesize that fine-tuning a foundation model to adapt to downstream tasks such as remote sensing segmentation primarily requires the adjustment of deeper semantic layers. Therefore, we initially inserted adapters into all four stages and gradually moved toward inserting them only in the deeper stages. Experimental results reveal that adding adapters to early transformer stages not only fails to improve domain adaptation, but can even hinder the model’s ability to specialize for remote sensing data. Conversely, limiting adapter placement to only the final stage also leads to underfitting, as the model lacks sufficient capacity for structural adaptation. We observe that inserting adapters into Stage 3 and Stage 4 yields the best performance, striking a balance between semantic depth and structural flexibility. While our experiments primarily report results using the proposed FreqWeaver Adapter, we also validated this stage-wise trend using other adapter designs, including the SAM Adapter and Earth Adapter, and found consistent behavior across all cases.
@@ 5. Conclusion
In this work, we proposed a parameter-efficient semi-supervised segmentation framework for high-resolution remote sensing imagery. By integrating a domain-adaptive FreqWeaver Adapter with the SAM2 backbone and employing an uncertainty-aware teacher student strategy, our method addresses the challenges of limited annotations, scale variation, and structural consistency. Experiments show that our approach achieves strong segmentation performance with minimal parameter overhead, outperforming existing PEFT and high-resolution segmentation methods. These results highlight the potential of combining spectral decomposition and semi-supervised learning for effective foundation model adaptation in remote sensing.

### v2

@@ 4.4. Comparison With State-of-the-Art PEFT Methods
To evaluate the effectiveness of the proposed adapter, we conduct extensive comparative experiments against several state-of-the-art parameter-efficient fine-tuning (PEFT) methods, including the SAM Adapter (Wu et al., 2025), LoRA (Hu et al., 2022), Conv-LoRA (Zhong et al., 2024), and the Earth Adapter (Hu et al., 2025). All implementations are based on SAM2, ensuring a consistent backbone for fair comparison. These approaches represent diverse strategies for adapting SAM2, with different focuses on domain specificity and parameter efficiency. The comparison results are shown in Table 1.
@@
Limitations of LoRA-Based Methods. Our experimental results show that both LoRA and Conv-LoRA achieve competitive performance with very low parameter overhead. LoRA leverages low-rank decomposition for efficient adaptation and is widely used in lightweight fine-tuning tasks. However, due to its linear structure, it has limited ability to capture the spatial complexity of high-resolution remote sensing scenes, which often contain multi-scale objects, intricate semantics, and fine-grained textures. To mitigate this limitation, Conv-LoRA extends the original LoRA design by incorporating convolutional layers to enhance local feature extraction. This modification improves its capacity to model small objects and boundary details, resulting in measurable performance gains. Nevertheless, despite these improvements, Conv-LoRA still falls short in fully exploiting the potential of high-resolution remote sensing imagery, particularly in capturing multi-scale spatial structures and fine-grained textures, which indicates the need for further refinement.
@@
Limitations of Adapter-Based Methods. The SAM Adapter is originally developed for medical image segmentation and shows a certain level of adaptability when transferred to remote sensing tasks. However, medical images usually contain relatively simple structures with clear boundaries, while remote sensing data have higher resolution, more complex scene variations, and richer semantic categories. As a result, the SAM Adapter has limited ability to model fine-grained structures and semantic differences, and its performance improvement remains modest. In contrast, the Earth Adapter is specifically designed for remote sensing tasks, but its attempt to decouple frequency components lacks targeted high- and low-frequency modeling mechanisms. This limitation prevents it from fully capturing complex spatial structures and texture features, leading to higher parameter overhead and unsatisfactory overall performance.
@@
Effectiveness of FreqWeaver Adapter. Our FreqWeaver Adapter introduces a frequency-aware design for remote sensing applications. It explicitly decouples low- and high-frequency components and applies targeted processing to each, which suppresses low-frequency artifacts while preserving high-frequency structural details. This improves segmentation accuracy and allows better adaptation to the SAM2 backbone. Compared with the best-performing Conv-LoRA, FreqWeaver increases IoU by 1.78% and Dice by 2.8%, showing clear gains in both overall segmentation accuracy and fine-grained detail modeling. In addition, FreqWeaver achieves this higher performance with fewer parameters than both the Earth Adapter and the SAM Adapter, demonstrating strong domain-specific effectiveness and efficiency for high-resolution remote sensing tasks. We also provide visualizations to further validate its effectiveness, as shown in Fig. 4.
@@
Table 2. Quantitative Comparison with state-of-the-art Land Use and Land Cover (LULC) methods.
@@
4.5. Comparison with State-of-the-Art LULC Methods
@@
To evaluate the effectiveness of our proposed approach, we conduct extensive comparative experiments against several representative state-of-the-art Land Cover and Land Use Classification (LULC) methods. Specifically, we select five recent and competitive baselines from the literature (2021–2024)—ABCNet (Li et al., 2021), UNetFormer (Wang et al., 2022), CMTFNet (Wu et al., 2023), RS3Mamba (Ma et al., 2024b), and SAM-RS (Ma et al., 2024a)—chosen for their strong benchmark performance and representation of diverse architectural paradigms, including CNN-based, Transformer-based, and Mamba-based models. The quantitative comparison results are shown in Table 2, while qualitative visualization examples are presented in Fig. 5 to further illustrate the advantages of our method.
@@
Limitations of Existing Methods. Quantitative results (Table 2) demonstrate that the proposed method consistently outperforms all selected baseline methods. In comparison to the strongest baseline, SAM-RS, our model achieves an improvement of 3.44% in IoU and 2.03% in Dice score, underscoring its superior performance in LULC tasks. Despite the reasonable segmentation accuracy exhibited by prior methods, visual analysis reveals several limitations: (1) Misclassification remains prevalent in complex scenes, particularly those with high intra-class variance or subtle inter-class distinctions (e.g., differentiating various types of vegetation or urban structures); (2) Object boundaries are often blurry and imprecise, a critical deficiency for ultra-high-resolution segmentation tasks that demand sharp delineations; (3) Models generally perform well on spectrally distinct classes like grass and trees, but falter on categories with complex geometries or semantic overlaps, such as distinguishing between parking lots and buildings, or between roads and barren land. Moreover, small object detection, such as identifying pools in residential areas, is often overlooked by existing approaches.
@@
Effectiveness of Our Method. We adapt the SAM2 foundation model for the LULC task, leveraging its rich prior knowledge to obtain robust category- and pixel-level representations. Building upon this foundation, our FreqWeaver Adapter enhances the semantic granularity of feature modeling by explicitly capturing and integrating multi-frequency information crucial for detail preservation, thereby enabling significantly improved delineation of object boundaries and detection of small structures. Notably, the enhanced feature representation from the FreqWeaver Adapter allows our model to effectively exploit contextual cues to better distinguish between semantically similar regions such as parking lots and roads. These capabilities demonstrate the strengths of our architecture in ultra-high-resolution LULC tasks. In addition, the integration of a semi-supervised learning strategy further improves the robustness and generalization of our model under limited annotation scenarios.
@@
Figure 6. Visualization of Land Cover Classification on the Baltimore City Map. The results show that our method achieves reliable land cover classification across the entire city at ultra-high spatial resolution, clearly revealing the overall urban structure.
4.6. Downstream Applications in Ultra-High Spatial Resolution Map Analysis
To evaluate the practicality of our approach in real-world scenarios, we apply the proposed method to downstream ultra-high-resolution map analysis tasks, as summarized in Table 3. Despite being trained on a relatively small dataset, our method exhibits strong generalization capabilities, achieving an Overall Accuracy (OA) of 84.55% and a Kappa coefficient of 81.38%. These metrics reflect not only the correctness of predictions but also the statistical reliability of classification performance across diverse land cover categories.
A detailed breakdown of class-wise performance is shown in Table 3. Notably, our model achieves high User’s Accuracy for categories such as Building (93.71%), Water (94.92%), and Grass/Shrub (91.80%), indicating precise identification of spatial regions with distinct spectral or structural characteristics. At the same time, it maintains balanced Producer’s Accuracy, particularly for Tree Canopy (98.98%) and Road (89.94%), demonstrating strong recall and comprehensive spatial coverage.
@@ 4.6. Downstream Applications in Ultra-High Spatial Resolution Map Analysis
To further validate the effectiveness of our method, we perform qualitative visualizations on real-world map segments, as illustrated in Fig. 6. The results show that our approach successfully preserves object boundaries, captures fine-grained structures, and adapts effectively to spatial heterogeneity, confirming its suitability for downstream high-resolution map analysis tasks.
4.7. Ablation Studies
To evaluate the contribution of each component within the proposed framework, we conduct extensive ablation studies, as shown in Table 4. In each experiment, specific modules—FreqWeaver (FW) and the Uncertainty-Aware Teacher–Student Framework (UATS)—are selectively enabled or disabled to analyze their individual impact on segmentation performance.
Effectiveness of FreqWeaver. We introduce FreqWeaver (FW) to decouple the frequency components and apply specific modeling within SAM2, enabling a parameter-efficient mechanism to capture complex patterns in remote sensing imagery. Experimental results show that, compared to full fine-tuning, our method achieves a 2.81% improvement in IoU and a 3.25% improvement in Dice score, while updating only 5.96% of the total parameters. These findings highlight the effectiveness of FreqWeaver in fine-tuning large vision models for remote sensing tasks and provide a scalable pathway for adapting foundation models under strict resource constraints.
@@ 4.8. Segment Anything 2 Adaptation
Segment Anything 2 is built upon the Hierarchical Detection Transformer (HieraDet), a multi-scale transformer architecture composed of four stages. In the original Segment Anything framework, adapters such as the SAM Adapter are inserted into every transformer block. However, directly applying this strategy to Segment Anything 2 often results in suboptimal fine-tuning performance due to its deeper and hierarchically structured design. To better accommodate these architectural differences, we conduct a series of empirical studies to investigate more effective adapter integration strategies. The results are shown in Table 5.
Specifically, we hypothesize that fine-tuning a foundation model to adapt to downstream tasks such as remote sensing segmentation primarily requires the adjustment of deeper semantic layers. Therefore, we initially insert adapters into all four stages and gradually move toward inserting them only in the deeper stages. Experimental results reveal that adding adapters to early transformer stages not only fails to improve domain adaptation but can even hinder the model’s ability to specialize for remote sensing data. Conversely, limiting adapter placement to only the final stage also leads to underfitting, as the model lacks sufficient capacity for structural adaptation. We observe that inserting adapters into Stage 3 and Stage 4 yields the best performance, striking a balance between semantic depth and structural flexibility. While our experiments primarily report results using the proposed FreqWeaver Adapter, we also validate this stage-wise trend using other adapter designs, including the SAM Adapter and Earth Adapter, and find consistent behavior across all cases.
@@ 5. Conclusion
In this work, we present Baltimore Atlas, a framework that combines the FreqWeaver Adapter and the Uncertainty-Aware Teacher Student Framework for ultra-high spatial resolution land cover classification. The FreqWeaver Adapter separates image information into high- and low-frequency components, modeling each to capture both global and local details, and adapts SAM2 to UHSR data with minimal parameters. The Uncertainty-Aware Teacher–Student Framework uses high-confidence pseudo-labels to learn more reliable features from unlabeled data. Combined, these components support fine-grained UHSR land cover classification and reduce the need for large labeled datasets. Experiments on the Baltimore Atlas dataset show that our FreqWeaver Adapter outperforms leading parameter-efficient tuning methods, and the full framework achieves better results than current state-of-the-art land cover classification approaches.